29. A mutual fund manager has a $40.00 million portfolio with a beta of 1.00. The risk-free rate is 4.25%, and the market risk premium is 6.00%. The manager expects to receive an additional $29.50 million which she plans to invest in additional stocks. After investing the additional funds, she wants the fund's required and expected return to be 13.00%. What must the average beta of the new stocks be to achieve the target required rate of return? Do not round your intermediate calculations.
Answer #1

Expected ret = Rf + Beta ( Market risk Premium )

13% = 4.25% + Portfolio Beta ( 6%)

Portfolio Beta = [ 13% - 4.25% ] / 6%

= 8.75% / 6%

= 1.46

Portfolio Beta is weighted Avg Beta of Securities in that Portfolio.

Particulars Weight Beta Wtd Beta
Portfolio 1 0.57554 1 0.5755
Additional Funds 0.42446 X 0.4245 X
Portfolio Beta 0.4245X + 0.5755

Thus 0.4245X + 0.5755 = 1.46

0.4245X = 1.46 - 0.5755

= 0.8845

X = 0.8845 / 0.4245

2.08

Beta of Additional funds is 2.08

option A is correct.